A certain weak base has a K, of 8.70 x 10-7. What concentration of this base will produce a pH of 10.147 concentration:
Answer #1

ko for the weak base given - Ko = 8.70 x 10-7 a pkb = – log ko = - log (8.70 810-7) = 6.06 Let, concentration of weak base bea / log c = 10.14 - 10.97 a 1/2 loge = – 0.83 a logc = -1.66 a c = 10 -1.66 = 0.0219 [three significant - figures] Thus, theI have given the answer in three significant figures......if the answer does not match with the computer database due to significant figures then please comment below.......